# Pipeline Parallelism with Controllable Memory

### v3

@@ Abstract
Pipeline parallelism has been widely explored, but most existing schedules lack a systematic methodology. In this paper, we propose a framework to decompose pipeline schedules as repeating a building block and we show that the lifespan of the building block decides the peak activation memory of the pipeline schedule. Guided by the observations, we find that almost all existing pipeline schedules, to the best of our knowledge, are memory inefficient. To address this, we introduce a family of memory efficient building blocks with controllable activation memory, which can reduce the peak activation memory to 1/2 of 1F1B without sacrificing efficiency, and even to 1/3 with comparable throughput. We can also achieve almost zero pipeline bubbles while maintaining the same activation memory as 1F1B. Our evaluations demonstrate that in pure pipeline parallelism settings, our methods outperform 1F1B by from 7% to 55% in terms of throughput. When employing a grid search over hybrid parallelism hyperparameters in practical scenarios, our proposed methods demonstrate a 16% throughput improvement over the 1F1B baseline for large language models.
@@ 1 Introduction
Although PP has been widely adopted and developed, it suffers from two prominent disadvantages: device idle time (referred to as pipeline bubbles) and large activation memory. To eliminate pipeline bubbles, one line of work focuses on asynchronous PP (Gaunt et al., 2017; Harlap et al., 2018; Yang et al., 2021) which is theoretically bubble free. However, it sacrifices the exact optimization semantics and may result in lower convergence performance (Lian et al., 2018; Tang et al., 2020). A parallel line of works revolve around synchronous PP, focusing on reducing pipeline bubbles and/or activation memory. GPipe (Huang et al., 2019) is a prominent early work to reduce the bubble rate by increasing the number of microbatches, at the cost of more activation memory. 1F1B (Fan et al., 2021) avoids the activation memory growth with respect to the number of microbatches by smartly staggering forward pass and backward pass, while keeping the same bubble rate with GPipe. Another notable work is GEMS (Jain et al., 2020), which stores activation memory of only one forward pass by scheduling microbatches one after another among two model replicas, thus with a significantly large bubble rate. Chimera (Li and Hoefler, 2021) extends the ideas of GEMS by combining two pipelines in different directions together, which reduces pipeline bubbles but with doubled parameter memory. In Megatron-LM (Narayanan et al., 2021), an interleaved strategy is proposed to further reduce the bubble rate, at the cost of more communication cost and a portion of extra activation memory. BPipe (Kim et al., 2023) focuses on reducing the activation memory of 1F1B from another perspective, transferring activation memory across devices based on the memory imbalance of 1F1B. However, it introduces a lot of extra communication and increases the complexity of the system, which makes it inefficient especially in settings with limited bandwidth. Zero Bubble (Qi et al., 2023) splits the backward into activation gradient computation and weight gradient computation, which can either reduce the pipeline bubbles without changing the maximum peak activation memory, or achieve zero bubble at the cost of doubled activation memory compared to 1F1B.
@@
In this paper, we first demonstrate all existing pipelines can be seen as repeating a basic building block in time. We then identify a direct link between the activation memory and the lifespan of each building block, which reveals the core insight of this paper: lifespan decides the activation memory. Based on this insight, we present a family of novel and memory-efficient building blocks and their pipelines. Compared to 1F1B, we reduce the activation memory to 1/2 asymptotically with even higher throughput, and to 1/3 asymptotically with comparable throughput. We can also achieves zero bubble under the same activation memory with 1F1B. Notably, our strategy is almost a pure gain to the existing methods, only at the cost of doubled communication cost between pipeline stages, which is relatively small and can be neglected.
@@ 2 How to Build a Pipeline
Building Block: It starts by laying out the passes for a single microbatch, which we call a building block. For example, the building block of 1F1B schedule is made of a sequence of forward passes followed by backward passes in the reverse order. We highlight the building block of 1F1B in color in Figure 1(a).
Repeating: More microbatches are then introduced. The building blocks are repeated and woven together to form a pipeline. In Figure 1 (top), the repeating building blocks are shown in different shades of grey. Notably, legit building blocks are required to repeat without a collision, namely, the passes from two building blocks should not overlap with each other.
@@
Figure 1: How to build a pipeline.
@@ 2.1 Building Blocks
The computation and memory efficiency of various pipelines can be attributed to their building blocks. The diversity of the building blocks primarily comes from three factors, model partitioning, device placement, and offsets between passes. We follow the idea and notations in zero bubble PP (Qi et al., 2023), using F to denote forward pass, B to denote “backward for the input”, and W to denote “backward for the weights”. Note that such finer granularity can be generalized to previous methods like 1F1B, by always grouping B and W together.
@@
Device placement is another key factor in the design of building blocks. While conventionally each pipeline stage is sequentially placed on a different device, it is not uncommon to place multiple stages on the same device like in interleaved 1F1B. Another example of unconventional device placement is Chimera (Li and Hoefler, 2021), where two pipelines are placed in reversed device order.
Last but not least, the offsets between F,B,W passes play a major role in the computation and memory efficiencies of the pipeline. By simply enlarging the offsets between subsequent F passes in the building block of 1F1B, we obtain the eager 1F1B (Zhuang et al., 2023) (Figure 1(b)) where more F passes are eagerly scheduled, resulting in higher memory consumption. GPipe (Huang et al., 2019) can be seen as adding a large offset between the last F and the first B in the 1F1B building block. One more example on the effect of the offset is in the comparison of ZB-H1 and ZB-H2 schedules, one can see that properly chosen offsets result in zero bubble schedules like ZB-H2. In this work, we assume that every F, B and W pass takes equally one unit of computation time, and only consider integer unit of offsets. Although this may limit the number of feasible building blocks, it greatly improves the simplicity of analysis.
@@ 2.2 Calculating the Peak Memory
Two quantities are crucial for the calculation of peak memory, the lifespan of a stage, and the repeating interval of the building blocks, both of which are illustrated in Figure 1. The lifespan of a stage is defined as the amount of time in between the starting of the F pass and the ending of B or W pass. A piece of activation memory is allocated at the starting of F, and retained in the RAM throughout the lifespan until it is consumed by both B and W. The peak memory consumption can be calculated by finding the maximum number of microbatches whose lifespans overlap with that of every other microbatch. Using $l$ to denote lifespan, $T$ to denote the repeating interval and $m$ the size of activation memory for a single microbatch, we have the relation.
@@ 2.3 Repeating without Collision
One constraint to keep in mind when designing the building blocks is that a legit building block is required to repeat without any collision. It may seem unintuitive how to design building blocks with this constraint. In practice, we design the building block first and perform a post-hoc verification. Another useful observation is that a legit building block usually produces a stable phase in the middle of pipeline, which contains a repeating $d\times T$ rectangle, where $d$ is the number of devices and $T$ is the repeating interval. This offers an alternative to constrain the building blocks. We can start by ordering passes within this rectangle and convert it back to a building block.
@@ 3 Memory Efficient Building Blocks
With the above framework, we can conveniently analyse the memory consumption pattern of existing pipelines. To our knowledge, all existing pipelines are memory inefficient due to two primary reasons: redundant dependency chain, and imbalanced memory usage. Before Zero Bubble (Qi et al., 2023), the backward is often regarded as a single pass, resulting in unnecessarily longer lifespan thus more memory footprint. In this paper, we leverage the backward splitting strategy to remove these redundant lifespan. The imbalanced memory comes from the innate heterogeneity of the lifespans across stages. From Figure 1(a), we can easily see that the lifespan of the stages differs greatly from each other, with the first stage having the longest lifespan. Consequently, it causes a memory bottleneck on the first device and under utilization of memory on all other devices. To resolve this problem, we introduce a family of novel building blocks which we refer to as V-Shape building blocks. The core insight comes from Equation 1 which says that the peak memory depends on the sum of the lifespans. Therefore, when we place multiple stages on the same device, we should always collocate stages of long lifespans with those of short lifespans. When the total sum of lifespans is fixed, balanced placement always means higher memory efficiency. This can be demonstrated by Figure 2, the parallel schedule (used in interleaved 1F1B) is imbalanced and has a memory bottleneck proportional to $l_{1}+l_{4}$, while in the V-Shape schedule it is $l_{1}+l_{6}$.
@@
Figure 2: Parallel vs V-Shape.
@@ 3.1 Controllable Balanced Memory
In this section, we introduce a simple approach to control the peak memory by using uniform offsets within each half of F and B passes. Note that we are aiming at the memory efficiency, so we are only interested in the building blocks with balanced peak memory across devices.
@@
Figure 3: V-Shape building blocks.
@@
Figure 4: V-Shape schedules compared to 1F1B.
We assume the model is uniformly partitioned, namely, both the computation and memory of each stage is the same. For a single microbatch, we denote the activation memory of each stage as $m$, and the total activation memory of the entire model as $M$. Note that $M=2dm$, where $d$ is the number of devices. To reduce the search space, we use uniform offsets within each half of F and B passes. Specifically, we apply the same offset (denoted as $\delta_{F}^{0}$) between two adjacent F passes within the first $d$ stages. For examples, we set $\delta_{F}^{0}$ as $2$ in Figure 3(b), as $1$ in Figure 3(c) and as $4$ in Figure 3(d). We add similar constraints for the other half of F passes and both half of B passes, and use $\delta_{F}^{1},\delta_{B}^{0},\delta_{B}^{1}$ to denote the offset between adjacent F passes within the last $d$ stages, the offset between adjacent B passes within the first $d$ stages, the offset between adjacent B passes within the last $d$ stages, respectively. To guarantee balanced peak memory across devices, we add another two constraints, $\delta_{F}^{0}=\delta_{B}^{1}=\delta^{0}$ and $\delta_{F}^{1}=\delta_{B}^{0}=\delta^{1}$, where we use notations $\delta^{0}$ and $\delta^{1}$ for simplicity. As in Figure 3(d), we set $\delta^{0}=\delta_{F}^{0}=\delta_{B}^{1}=4$ and $\delta^{1}=\delta_{F}^{1}=\delta_{B}^{0}=2$. Note that we only control the offsets across different devices. For those adjacent passes within the same device, e.g. F and B of the last stage, B and W for all stages, we leave them flexible and only ensure their offsets are small (less than the repeating interval). According to Equation 1, we can analyse the asymptotic peak memory with respect to $d$,
@@ 3.2 V-Shape Pipeline Schedules
By varying the values of $\delta^{0}$ and $\delta^{1}$, we come up with 3 novel V-Shape building blocks (Figure 3), and present their final schedules based on our framework in Figure 4. The building block of V-Min (Figure 3(c)) has the minimum offsets, namely $\delta^{0}=\delta^{1}=1$, thus the minimum memory consumption. With $\delta^{0}=4$ and $\delta^{1}=2$ as in Figure 3(d), V-ZB eliminates the bubble to almost 0 (Figure 4(d)), pushing to extreme throughput. The building block of V-Half (Figure 3(b)), which uses $\delta^{0}=2$ and $\delta^{1}=1$, sits between the two extremes and consumes half of the activation memory required by 1F1B. Although both V-Min and V-Half have lower memory footprint than 1F1B, V-Min contains about 2/3 and V-Half contains about 1/2 of 1F1B’s bubbles, assuming F, B, W have equal run time. We show the comparison between our proposed V-Shape schedules and 1F1B in Table 1.
Table 1: For peak memory and bubbles, we ignore those small constant values. We redefine $\delta^{0}$/$\delta^{1}$ as the offsets between adjacent forward/backward passes for 1F1B.
Building Block $\delta^{0}$ $\delta^{1}$ Peak Memory Bubbles
1F1B $2$ $4$ $M$ $6d$
V-Min $1$ $1$ $M/3$ $4d$
V-Half $2$ $1$ $M/2$ $3d$
V-ZB $4$ $2$ $M$ 00
Notably, to avoid collision in building block, the offsets for V-Min and V-Half is slightly different on different $d$. The details are in Appendix F.
@@ 3.3 Repeating Bubbles in V-Min
In real cases where F, B and W have different run times, V-Min suffers from a repeating bubble. As shown in Figure 5, there exists a bubble for every repeating interval $T$. As a result, the bubble grows with respect to the number of microbatches. Although V-Half may also encounter the same issue, we surprisingly find that it generates patterns that tessellates well in most empirical cases. As illustrated in Figure 5(b), the throughput of V-Half is robust to the variation of run times. Additionally, we show that the bubbles of V-ZB will never grow when increasing the number of microbatches. We leave the related discussions in Appendix E.

### v4

@@ Abstract
Pipeline parallelism has been widely explored, but most existing schedules lack a systematic methodology. In this paper, we propose a framework to decompose pipeline schedules as repeating a building block, and show that the lifespan of the building block decides the peak activation memory of the pipeline schedule. Guided by the observations, we find that almost all existing pipeline schedules, to the best of our knowledge, are memory inefficient. To address this, we introduce a family of memory efficient building blocks with controllable activation memory, which can reduce the peak activation memory to 1/2 of 1F1B without sacrificing efficiency, and even to 1/3 with comparable throughput. We can also achieve almost zero pipeline bubbles while maintaining the same activation memory as 1F1B. Our evaluations demonstrate that in pure pipeline parallelism settings, our methods outperform 1F1B by from 7% to 55% in terms of throughput. When employing a grid search over hybrid parallelism hyperparameters in practical scenarios, our methods demonstrate a 16% throughput improvement over the 1F1B baseline for large language models. The implementation is open-sourced at this url.
@@ 1 Introduction
Although PP has been widely adopted and developed, it suffers from two prominent disadvantages: pipeline bubbles and large activation memory. To eliminate pipeline bubbles, one line of work focuses on asynchronous PP (Gaunt et al., 2017; Yang et al., 2021), which is theoretically bubble free. However, it sacrifices the exact optimization semantics and may result in lower convergence performance (Lian et al., 2018; Tang et al., 2020). A parallel line of works revolve around synchronous PP, focusing on reducing pipeline bubbles and/or activation memory. GPipe (Huang et al., 2019) is an early work to reduce the bubble rate by increasing the number of microbatches, at the cost of more activation memory. 1F1B (Fan et al., 2021) avoids the activation memory growth with respect to the number of microbatches by staggering forward pass and backward pass, keeping the same bubble rate with GPipe. Another notable work is GEMS (Jain et al., 2020), which stores activation memory of only one forward pass by scheduling microbatches one after another among two model replicas, thus with a significantly large bubble rate. Chimera (Li and Hoefler, 2021) extends the ideas of GEMS by combining two pipelines in different directions together, which reduces pipeline bubbles when the number of microbatches is small, but with doubled parameter memory. Hanayo (Liu et al., 2023) is introduced to attain the same scheduling efficiency with Chimera without replicated models, but still suffering from scaling to more microbatches. Although its wave-like scheme is kind of similar to our V-shape building blocks, it is not motivated for memory balance, thus resulting in totally different pipeline schedules. In Megatron-LM (Narayanan et al., 2021), an interleaved strategy is proposed to further reduce the bubble rate, at the cost of more communication cost and a portion of extra activation memory. BPipe (Kim et al., 2023) focuses on reducing the activation memory of 1F1B from another perspective, transferring activations across devices based on the memory imbalance of 1F1B. However, it introduces a lot of extra communication and increases the complexity of the system, which makes it inefficient especially in settings with limited bandwidth. Zero Bubble (Qi et al., 2023) splits the backward into activation gradient computation and weight gradient computation, which can either reduce the pipeline bubbles without changing the maximum peak activation memory, or achieve zero bubble at the cost of doubled activation memory compared to 1F1B.
@@
In this paper, we first demonstrate all existing pipelines can be seen as repeating a basic building block in time. We then identify a direct link between the activation memory and the lifespan of each building block, which reveals the core insight of this paper: lifespan decides the activation memory. Based on this insight, we present a family of novel and memory-efficient building blocks and their pipelines. Compared to 1F1B, we reduce the activation memory to 1/2 asymptotically with even higher throughput, and to 1/3 asymptotically with comparable throughput. We can also achieve zero bubble under the same activation memory with 1F1B. Notably, our strategy is almost a pure gain to the existing methods, only at the cost of doubled communication cost between pipeline stages, which is relatively small and can be neglected.
@@ 2 How to Build a Pipeline
Building Block: It starts by laying out the passes for a single microbatch, which we call a building block. For example, the building block of 1F1B is made of a sequence of forward passes followed by backward passes in the reverse order. We highlight the building block of 1F1B in color in Figure 1(a).
Repeating: More microbatches are then introduced. The building blocks are repeated and woven together to form a pipeline. In Figure 1 (top), the repeating building blocks are shown in different shades of gray. Notably, legit building blocks are required to repeat without a collision, namely, the passes from two building blocks should not overlap with each other.
@@
Figure 1: A pipeline can be built by repeating a building block, and then squeezing redundant bubbles.
@@ 2.1 Building Blocks
The computation and memory efficiency of various pipelines can be attributed to their building blocks. The diversity of the building blocks primarily comes from three factors, model partitioning, device placement, and offsets between passes. We follow the idea and notations in zero bubble PP (Qi et al., 2023), using F to denote forward pass, B to denote “backward for the activations”, and W to denote “backward for the weights”. Note that such finer granularity can be generalized to previous methods like 1F1B, by always grouping B and W together.
@@
Device placement is another key factor in the design of building blocks. While conventionally each pipeline stage is sequentially placed on a different device, it is not uncommon to place multiple stages on the same device like in interleaved 1F1B (Figure 18(h)). Another example of unconventional device placement is Chimera, where two pipelines are placed in reversed device order.
Last but not least, the offsets between F,B,W passes play a major role in the computation and memory efficiencies of the pipeline. By simply enlarging the offsets between subsequent F passes in the building block of 1F1B, we obtain the eager 1F1B (Zhuang et al., 2023) (Figure 1(b)) where more F passes are eagerly scheduled, resulting in higher memory consumption (but better communication overlapping). GPipe can be seen as adding a large offset between the last F and the first B in the 1F1B building block. One more example on the effect of the offset is the comparison of ZB-H1 (Figure 18(c)) and ZB-H2 (Figure 18(d)) schedules, one can see that properly chosen offsets result in zero bubble schedules like ZB-H2. In this work, we assume that every F, B or W pass takes equally one unit of computation time, and only consider integer unit of offsets. Although this may limit the number of feasible building blocks, it greatly improves the simplicity of analysis.
@@ 2.2 Calculating the Peak Memory
Two quantities are crucial for the calculation of peak memory, the lifespan of a stage, and the repeating interval of the building blocks, both of which are illustrated in Figure 1. The lifespan of a stage is defined as the amount of time between the starting of the F pass and the ending of B or W pass. A piece of activation memory is allocated at the starting of F, and retained in the RAM throughout the lifespan until it is consumed by both B and W. The peak memory consumption can be calculated by finding the maximum number of microbatches whose lifespans overlap with that of every other microbatch. Using $l$ to denote lifespan, $T$ to denote the repeating interval and $m$ the size of activation memory for a single microbatch, we have the relation.
@@ 2.3 Repeating without Collision
One constraint to keep in mind when designing the building blocks is that a legit building block is required to repeat without any collision. It may seem unintuitive how to design building blocks with this constraint. In practice, we design the building block first and perform a post-hoc verification. Another useful observation is that a legit building block usually produces a stable phase in the middle of the pipeline, which contains a repeating $d\times T$ rectangle, where $d$ is the number of devices and $T$ is the repeating interval. This offers an alternative to constrain the building blocks. We can start by ordering passes within this rectangle and convert it back to a building block.
@@ 3 Memory Efficient Building Blocks
With the above framework, we can conveniently analyze the memory consumption pattern of existing pipelines. To our knowledge, all existing pipelines are memory inefficient due to two primary reasons: redundant dependency chain, and imbalanced memory usage. Before Zero Bubble (Qi et al., 2023), the backward is often regarded as a single pass, resulting in unnecessarily longer lifespan thus more memory footprint. In this paper, we leverage the backward splitting strategy to remove these redundant lifespan. The imbalanced memory comes from the innate heterogeneity of the lifespans across stages. From Figure 1(a), we can easily see that the lifespan of the stages differs greatly from each other, with the first stage having the longest lifespan. Consequently, it causes a memory bottleneck on the first device and under utilization of memory on all other devices. To resolve this problem, we introduce a family of novel building blocks, which we refer to as V-Shape building blocks. The core insight comes from Equation 1 which says that the peak memory depends on the sum of the lifespans. Therefore, when we place multiple stages on the same device, we should always collocate stages of long lifespans with those of short lifespans. When the total sum of lifespans is fixed, balanced placement always means higher memory efficiency. This can be demonstrated by Figure 2, the parallel schedule (used in interleaved 1F1B) is imbalanced and has a memory bottleneck proportional to $l_{1}+l_{4}$, while in the V-Shape schedule it is $l_{1}+l_{6}$.
@@
Figure 2: The V-Shape building block ensures balanced peak memory across all devices, whereas the parallel building block has a memory bottleneck in the first device.
@@ 3.1 Controllable Balanced Memory
Figure 3: V-Shape building blocks with 4 devices ($d=4$), where white text colors represent the first half of model stages and black text colors represent the second half. F, B and W represent the forward, backward (for activation gradients) and backward for weight gradients, respectively.
@@
Figure 4: V-Shape schedules compared to 1F1B, under the setting of 4 devices and 8 microbatches. The stable phases adhere to the pattern of their building blocks.
We assume the model is uniformly partitioned, namely, both the computation and memory of each stage are identical. For a single microbatch, we denote the activation memory of each stage as $m$, and the total activation memory of the entire model as $M$. Note that $M=2dm$, where $d$ is the number of devices. To make it simple and tractable, we use uniform offsets within each half of F and B passes to control the peak memory. Specifically, we apply the same offset $\delta_{F}^{0}$ between two adjacent F passes within the first $d$ stages (e.g., $\delta_{F}^{0}=2$ in Figure 3(b), $\delta_{F}^{0}=1$ in Figure 3(c) and $\delta_{F}^{0}=4$ in Figure 3(d)). Similar constraints are applied to the other half of the F passes and both halves of the B passes, denoted as $\delta_{F}^{1},\delta_{B}^{0},\delta_{B}^{1}$, respectively. To guarantee balanced peak memory across devices, we add another two constraints, $\delta_{F}^{0}=\delta_{B}^{1}=\delta^{0}$ and $\delta_{F}^{1}=\delta_{B}^{0}=\delta^{1}$, where we use notations $\delta^{0}$ and $\delta^{1}$ for simplicity. For example, in Figure 3(d), we set $\delta^{0}=4$ and $\delta^{1}=2$. Note that we only control the offsets across different devices. For those adjacent passes within the same device (e.g., F and B of the last stage, two F and two B in the last device), we use brute force to find optimal solutions, ensuring their offsets are small (less than the repeating interval). Note that W can always be placed greedily after settling all F and B passes, so we don’t need to search their offsets during brute force. According to Equation 1, we can analyze the asymptotic peak memory with respect to $d$,
@@
Table 1: Small constant values are ignored for bubbles and peak memory of V-Min and V-Half. For 1F1B, $\delta^{0}$/$\delta^{1}$ are redefined as the offsets between adjacent forward/backward passes. $M$ represents the total activation memory of the entire model, and $d$ is the number of devices.
Building Block $\delta^{0}$ $\delta^{1}$ Peak Memory Bubbles
1F1B $2$ $4$ $M$ $\approx 6d$
V-Min $1$ $1$ $\approx M/3$ $\approx 4d$
V-Half $2$ $1$ $\approx M/2$ $\approx 3d$
V-ZB $4$ $2$ $M$ 00
@@ 3.2 V-Shape Pipeline Schedules
By varying the values of $\delta^{0}$ and $\delta^{1}$, we come up with 3 novel V-Shape building blocks (Figure 3), and present their final schedules based on our framework in Figure 4. The building block of V-Min (Figure 3(c)) has the minimum offsets, namely $\delta^{0}=\delta^{1}=1$, thus the minimum memory consumption. With $\delta^{0}=4$ and $\delta^{1}=2$ as in Figure 3(d), V-ZB eliminates the bubble to almost 0 (Figure 4(d)), pushing to extreme throughput. The building block of V-Half (Figure 3(b)), which uses $\delta^{0}=2$ and $\delta^{1}=1$, sits between the two extremes and consumes about half of the activation memory required by 1F1B. Although both V-Min and V-Half have lower memory footprint than 1F1B, V-Min contains about 2/3 and V-Half contains about 1/2 of 1F1B’s bubbles, assuming F, B, W have equal run time. We show the comparison between our proposed V-Shape schedules and 1F1B in Table 1. Notably, the exact peak memory is $\lceil\frac{d+2}{3}\rceil\frac{M}{d}$ for V-Min, and $\lceil\frac{d+1}{2}\rceil\frac{M}{d}$ for V-Half. To avoid collisions in the building blocks of V-Min and V-Half, the offsets (within the same device) are slightly different for different values of $d$. The details are in Appendix F.
@@ 3.3 Repeating Bubbles in V-Min
In real-world scenarios where F, B and W have different run times, V-Min suffers from a repeating bubble. As shown in Figure 5, there exists bubbles for every repeating interval $T$. Consequently, the bubble grows as the number of microbatches increases. Although V-Half may encounter the same issue (when the times of F, B and W differ significantly), it generates patterns that tessellate well in most empirical cases due to its loose dependencies. As illustrated in Figure 5(b), the throughput of V-Half is robust to the variation of run times. Additionally, the bubbles of V-ZB will never grow when increasing the number of microbatches. We leave the related discussions in Appendix E.